OpenStudy (anonymous):

Angles θ and Φ are complementary, and θ =pi/12 Find the measure of Φ.????

OpenStudy (anonymous):

they add to \(\frac{\pi}{2}\) so \[\phi +\frac{\pi}{12}=\frac{\pi}{2}\] making \[\phi =\frac{\pi}{2}-\frac{\pi}{12}\]

OpenStudy (anonymous):

subtract as if the \(\pi\) wasn't there \[\frac{\pi}{2}-\frac{\pi}{12}=\frac{6\pi}{12}-\frac{\pi}{12}=\frac{5\pi}{12}\]